Three men, each having denarii, found a purse containing 23 denarii. The first
man said to the second, “If I take this purse, I will have twice as much as you.”
The second said to the third, “If I take this purse, I will have three times as much as you.” The third man said to the first, “If I take this purse, I will have four times as much as you.” How many denarii did each man have?

Answer by fractalier(6550)

Call what the three men had, a, b, and c.
Then we can write
a + 23 = 2b
b + 23 = 3c
c + 23 = 4a
We will take three equations and three unknowns and boil it down to two equations and two unknowns...
The first one becomes
(1/2)a + 11.5 = b
Plug that in to the second and get
(1/2)a + 11.5 + 23 = 3c
(1/2)a + 34.5 = 3c
or
(1/2)a - 3c = -34.5
or
a - 6c = -69
Let's combine this with the 3rd equation
-4a + c = -23
-24a + 6c = -138
Now add them
a - 6c = -69
+(-24a + 6c = -138)
--------------------
-23a = -207
a = 9
giving us
b = 16
and
c = 13